An intensity of a sound wave is measured to be 1.87×104 W/m2 at position A, 120 m away from the source of this sound. A worker at position B, 308 m away from the source hears the sound.

What is the energy of the sound wave at position B, if that sound lasts 5 minutes?

Select one:

1.0×1012 J

 1.69×1010 J 

 2.8×108 J

 8.4×109 J


1
Expert's answer
2021-11-21T17:30:57-0500

"I_2=I_1R_1^2/R_2^2=1.87\cdot10^4\cdot120^2/308^2=2838.59\ (W/m^2)"


"I=\frac{E}{A\cdot t}\to E=I\cdot A\cdot t=2838.59\cdot1\cdot5\cdot60=8.5\cdot10^5\ (J)"